Suppose that the weight of Florida navel oranges is normally distributed with mean µ = 8 ounces, and standard deviation σ = 1.5 ounces. (a) (1 point) State the model in notation form. (b) (2 points) What proportion of oranges weigh more than 11.5 ounces? (c) (2 points) What proportion of oranges weigh less than 8.7 ounces? (d) (2 points) What proportion of oranges weigh between 6.2 and 7 ounces? Page 3 (e) (5 points) What are the median, mode, first quartile, third quartile and IQR for the oranges? (f) (2 points) Find the 80th percentile of the distribution of X. (g) Suppose another strain of navel oranges from California has a mean µ = 8.5 ounces, and standard deviation σ = 1.7 ounces. A Florida orange is randomly selected and weighs 12 ounces, while a California orange is randomly selected and weighs 13 ounces. Which one strain of orange is ”relatively” bigger for its average size?
Answer #1

Given that M-8,0-1.15 a) Xe N( 8 1.5) b) P(x211.5) - 1-P (X= 11.5) = 1-PC Videos 11.58 1.5 1-PC2% 2.333) 21-0.9902 -0.0098 Poe) Median, Mean and Mode are equal for Normal Sandom Variable :: Median = 8 Modez8 QI => P(X+QI) -0.25 PC key < 0.53 )=0-25 PMr. Average weight of Californial naved oranges 80 percontic -e Po.80 P(XSP080)-0.80 P( 1-4The8 ) =0-80 plzs 20.80)-0.80 PC ZRR Console > #b) > pnorm (2.333) [1] 0.9901759 > #c) > pnorm (0.467) [1] 0.6797501 > #d) > pnorm (-0.67) [1] 0.2514289 [1] 0.